OpenStudy (anonymous):

∠1 and ∠2 are supplementary. m∠1 = 3x – 4 and m∠2 = 8x – 14. Find the measure of each angle.

OpenStudy (anonymous):

@Directrix

OpenStudy (mathstudent55):

Do you know what supplementary angles are?

OpenStudy (anonymous):

Angles that equal 180 degrees!

OpenStudy (anonymous):

Would i do this? (3x – 4)+(8x – 14)= 180

OpenStudy (mathstudent55):

Yes, 2 angles whose measures add up to 180 degrees. Angles 1 and 2 are supplementary. That means m<1 + m<2 = 180. Now substitute the expression for each angle and solve for x.

OpenStudy (mathstudent55):

Yes, your equation is correct. Now solve it for x. Then when you know x, plug it in the expression for each angle to find the measure of each angle.

OpenStudy (anonymous):

(11x – 18)= 180 That?

OpenStudy (mathstudent55):

Correct. Drop the parentheses. You don't need them.

OpenStudy (anonymous):

x= 18

OpenStudy (anonymous):

i think i got it! m∠1 = 50°, m∠2 = 130°

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

Great. Notice that 50 deg and 130 deg are the measures of supplementary angles, which makes sense.

OpenStudy (mathstudent55):

You're welcome.